OpenStudy (cenaida):

For the proportional relationship y = 2x, what is the ratio of vertical change to horizontal change? A. 2/1 The graph increases from left to right. B. 2/0 The graph increases from left to right. C. 2/1 The graph decreases from left to right. D. 2/0 The graph decreases from left to right.

OpenStudy (campbell_st):

slope as a proportion is rise/run so write the slope of the line 2... as a fraction.... 2/1 so \[\frac{rise}{run} = \frac{2}{1}\] positive slopes go from lower left to upper right.

OpenStudy (cenaida):

is it C then?

OpenStudy (campbell_st):

here is a line with a positive slope 

OpenStudy (cenaida):

then it is c? it decreases?